# Unpacking DPO and PPO: Disentangling Best Practices for Learning from Preference Feedback

### v1

@@ 1 Introduction
  • Increasing reward model size and dataset size used to train the reward model results in improved reward model performance on benchmarks directly testing reward model performance. Examining their effect on policy model performance when used during PPO training, we find that these improved reward models have a large effect on GSM performance, but give marginal to no improvements across all other evaluations considered (§3).
@@ 3.1 Preference Data
Preference-based learning has the strongest effect on instruction following and truthfulness performance.
Our best models improve on the SFT model by over 8 points in these categories. In contrast, preference-based learning does not aid factuality, with all models remaining with 1 point of each other. This suggests preference-based learning is most useful for improving chat-related abilities (instruction following, truthfulness) and learning stylistic features, but less strong at teaching new facts to a model. Interestingly, we observe that training on the Chatbot Arena data performs poorly on safety, suggesting Chatbot Arena volunteers generally prefer more toxic completions.
@@
Synthetic datasets generally outperform human-annotated and web-scraped datasets, especially in truthfulness. Additionally, using datasets collected by first getting per-aspect annotations (i.e., annotations from a human or model that score the helpfulness, harmlessness, etc. of the response independently) and then averaging across these scores tend to outperform datasets that rely only on overall judgements (i.e., just asking the annotator for an overall score instead of a per-aspect score). The two datasets that use this method, HelpSteer and UltraFeedback, display stronger or similar performance to datasets up to 15 times larger (e.g. HelpSteer vs HH-RLHF). We investigate the performance of varied sub-splits of UltraFeedback in App. E and find that the quality of pairwise judgements is highly important, rather than the quality of the underlying models used to generate the dataset.
@@
We now compare algorithms for learning from preferences, comparing the performance of DPO and PPO when the same base models and data are used (Table 2). We use exactly the same data for training DPO and PPO models, [^5] That is, we compare DPO models trained directly on each dataset with PPO models trained using reward models trained directly on that dataset and using prompts from the same dataset. Note that PPO uses additional generations from the model during training, but both approaches use the same amount of labelled data. and subsample larger datasets to 60,908 examples due to computational resources and only use these subsampled datasets during reward model and PPO training. Results in Table 2 indicate that:
@@
Across all datasets, models trained with PPO outperform models trained with DPO. In fact, PPO is able to provide improvements over the SFT model in cases where DPO training did not, such as when using StackExchange. On average, PPO improves over DPO performance by 0.7 points.
PPO improves reasoning, coding and safety capabilities most.
@@ 3.3 Reward Models
Results in Table 3 indicate that both increasing the reward model dataset (‘Mix’) and reward model size (from 13B to 70B) improve performance. Surprisingly, we find that our 70B reward models perform best on BoN evaluation, while the 13B mixture model performs best on RewardBench. Both evaluations show that increasing the dataset mixture and increasing the dataset size can help, although increasing the dataset mixture is not as useful for further improving the 70B reward model. Additionally, we note that our best performing reward model, the 13B mix RM, outperforms all other reward models based on Llama 2 (including 70B reward models) on RewardBench at time of writing.
@@
We then test if our improved reward models lead to improved downstream policy models when used during PPO training. We perform PPO training using the UltraFeedback prompts during training and report our results on the right side of Table 3. Surprisingly, we find that improvements in reward models result in marginal to no improvements in overall downstream performance. We see only small improvements in overall performance from using the 70B UltraFeedback RM, despite the fact that the RM performed significantly better than the 13B UltraFeedback RM on the RewardBench and Best-of-N evaluations. This improvement is largely driven by a large performance jump in GSM, as shown in Table 3, while other metrics stay largely similar. This suggests that it is difficult to translate improvements in reward models to the underlying policy. We note that most prior work examining reward models tends to examine either direct reward model performance [26, 61] or proxy reward [17, 40], rather than downstream performance. Our findings suggest that improving such evaluations may not necessarily translate to improved downstream performance. Additionally, we find that using the larger 70B UltraF. RM is less sensitive and performs well with lower KL penalty coefficient values than using the 13B UltraF. RM, and further examine the effect of the KL penalty coefficient on performance in App. I.
@@ 3.4 Policy Training Prompts
We first examine the effect of using policy training prompts targeted towards a particular downstream setting—specifically, math capabilities as evaluated by GSM. We construct three different prompt sets: prompts sourced from the GSM train set, math-related prompts from varied datasets, and 20K random prompts from UltraFeedback. We further detail the approach used to identify math-related prompts in Appendix J. We then train models using PPO with each reward model from the previous section on each prompt set. Results in Fig. 3 demonstrate that:
@@
Surprisingly, we observe that using mixed prompts does not seem to improve performance in the generalist setting. When looking at code and math results specifically, we do not see consistent improvements using the altered prompt mixture, and our overall performance tends to drop. This is likely due to the already diverse nature of UltraFeedback, such that when looking at the whole dataset (i.e., not just the 20K subset in Figure 3), we are able to reach strong performance on math and coding evaluations. Altering the distribution away from other tasks slightly hurts the overall performance.
@@ 4 A Recipe for Learning from Preferences
Putting together all our findings from previous sections, we suggest a recipe for training a strong model using learning from preferences, as shown in Figure 1 and in Table 5. We take a high-quality, synthetic preference dataset, a large reward model, [^7] However, a smaller 13B RM performs almost as well if one is compute-constrained. and train it using PPO. If we additionally wish to focus on a specific domain, we can additionally collect domain-specific prompts for policy training. We find that our best model (Tülu 2+PPO 13B trained with the 70B UltraF. RM) outperforms our best DPO model and other popular models based off Llama 2 13B, including Llama 2 Chat, which has also undergone SFT and PPO training. Additionally, incorporating task-specific prompts into policy training may further improve performance when the prompts align closely with downstream evaluations, as shown in Figure 3. Finally, we note that using a stronger base policy than Tülu 2 13B may also yield further improvements, although how to improve the pretraining and instruction-tuning phases is not the focus of this work.
Model Factuality Reasoning Coding Truthfulness Safety Instr. Foll. Average Llama 2 Chat 13B [52] 53.2 24.7 36.9 88.0 91.9 51.2 57.7 Nous Hermes 13B [51] 53.2 43.5 47.7 80.5 43.9 38.7 51.3 Vicuna 1.5 13B [64] 54.5 39.3 38.5 62.8 92.4 45.8 55.6 Llama 2 13B Base 52.0 37.0 30.7 32.7 32.7 - - Tülu 2 13B SFT 55.4 47.8 45.1 56.6 91.8 44.2 56.8 Tülu 2+DPO 13B 55.3 50.9 45.9 69.3 91.9 52.8 61.0 Tülu 2+PPO 13B (13B UFRM) 56.0 52.0 47.7 71.5 91.8 54.4 62.2 Tülu 2+PPO 13B (70B UFRM) 55.4 53.9 47.3 72.3 91.9 55.8 62.8 Tülu 2+PPO 13B (70B UFRM+MP) 55.3 53.1 48.4 71.0 92.7 54.0 62.4
Table 5: Putting together a recipe for preference-based learning: Performance of our best-performing models along with popular open models based on Llama 2 13B. ‘MP’ refers to using the mixed prompt set described in §4. Using PPO with a large reward model performs best overall.
@@ 6 Conclusion
In this work, we have systematically explored the core components of learning from preference feedback and examined the relative impacts of each in turn on model performance across a wide range of evaluations. Our results suggest a clear ordering of importance: preference data quality, algorithm choice, reward model quality, and finally targeted policy training prompts. Additionally, we find that using larger reward models can significantly improve math capabilities, but have marginal effects on other capabilities we evaluate in this work. Overall, we suggest a recipe for learning from preference feedback with currently available resources: best performance can be achieved by using a strong synthetic dataset (UltraFeedback), and using PPO training with a large reward model. Our work suggests that further exploring how to make better use of improved reward models is an important direction for further improving PPO-style approaches to learning from preference data. We plan to release models and code related to this paper and hope that our settings provide a firm base for future work further exploring learning from preferences for language models.
@@ Appendix A Limitations & Broader Impacts
As an empirical study with limited compute, our results are largely based on an in-depth examination over a single model suite (Tülu 2). As such, results may vary somewhat when applied to vastly different model families. Additionally, our evaluation may not reflect all potential downstream use-cases - for example, we do not test multilingual performance. As such, examining how our results carry to multilingual settings or greatly differing base models is interesting future work. However, we have done our best to explore a wide variety of evaluations and datasets, using publicly available resources to aid in reproducibility. In terms of computational efficiency, we note that, while PPO does well, it comes with a significantly increased computational cost compared to DPO. We only consider performance in this study and don’t explicitly measure the relative computational cost of different methods (e.g., by measuring FLOPs cost of training with DPO vs PPO). There are also other aspects to keep improving PPO (e.g., doing new rounds of preference annotations to update the reward model for the changing distributions of the policy model), which we cannot exhaust in this paper, and we leave the exploration of them for future work.
@@ Appendix B Dataset Details
We additionally construct our ‘mix’ used for expanding the reward model training by a number of examples from each dataset as listed in Table 6.
@@ Appendix E UltraFeedback Subset Study
Factuality Reasoning Coding Truthfulness Safety Inst. Foll. Average Llama 2 base 52.0 37.0 30.7 32.7 32.7 - - Tulu 2 (SFT) 55.4 47.8 45.1 56.6 91.8 44.2 56.8 UltraFeedback (FG; All) 55.3 50.9 45.9 69.3 91.9 52.8 61.0 Strong Model Gen.s 55.3 46.5 45.6 58.9 90.1 47.6 57.3 Middle Model Gen.s 55.2 48.5 44.5 60.6 91.5 48.2 58.1 Weak Model Gen.s 55.2 47.3 45.4 60.5 90.4 46.2 57.5 FalseQA Prompts 55.2 49.2 44.0 60.8 92.2 45.5 57.8 Evol Instruct Prompts 55.5 47.8 47.5 58.4 90.6 48.7 58.1 TruthfulQA Prompts 55.4 47.6 44.5 55.7 91.6 45.4 56.7 Flan V2 Prompts 55.3 49.4 46.3 62.4 91.3 48.9 58.9 ShareGPT Prompts 55.4 49.8 43.9 60.2 87.5 50.3 57.9 UltraChat Prompts 55.4 50.6 47.5 59.9 89.9 48.1 58.6
Table 7: Performance of varied sub splits of UltraFeedback, considering both samples using generations from models of varying strength (top, middle, strong model gen.) strong and only considering samples using prompts from single datasets (e.g., FalseQA, Flan V2). Sampling from different-quality models makes relatively little difference, while prompt choice matters. Additionally, different source datasets provide improvements in different evaluations.
@@ F.1 DPO
We found that the hyperparameters suggested by Tunstall et al. [53] and followed by Ivison et al. [22] worked well across varied datasets in pilot experiments. This involves 3 epochs of training with a learning rate of $5\times 10^{-7}$, with a linear warmup for 10% of training and linear cooldown for the duration. We use the open-sourced codebase used for training Tülu 2+DPO. [^14] https://github.com/hamishivi/EasyLM
@@ F.2 PPO
PPO formulates language generation as a Markov decision process (MDP), where each response $y$ is an episode, an action is a token $y_{t}$, and a state is the concatenation of a prompt and a partial response ($x\circ y_{<t}$). To construct token-level rewards $r_{t}$ that guide the RL training, the sequence-level reward $r$ is applied only to the last token, and each token is penalized by a KL term $-\beta\big{(}\log\pi_{\theta}(y_{t}\mid x\circ y_{<t})-\log\pi_{\text{ref}}(y_{t}\mid x\circ y_{<t})\big{)}$. Formally, the token-level reward $r_{t}$ for each response token $y_{t}$ is defined as
@@
The prompt and continuation each has at most 1024 tokens. We use a batch size of 64 and the same for minibatch size, and train for 1 epoch over the prompt dataset. On each batch of prompts and rollouts, we train for 1 inner epoch. We use a sampling temperature $\tau=0.7$ for rollout, and a fixed KL penalty coefficient $\beta=0.05$. We optimize with AdamW with learning rate $\eta=1\times 10^{-6}$ with a linear warmup for 10% of iterations. We provide additional details on the hyperparameters used for PPO in Table 9. In experiments, we found that training with larger (70B) reward models benefited from a reduced KL penalty coefficient $\beta$, but using smaller reward models did not benefit (see App I for further discussion). As such, for runs using larger reward models, we use a KL penalty of $\beta=0.0325$ instead of $\beta=0.05$. This is in line with prior work suggesting that improved reward models are more difficult to overoptimize against, and so can benefit from lower KL penalties [17, 40].

### v2

@@ 1 Introduction
  • Increasing reward model size or dataset size used to train the reward model results in improved reward model performance on benchmarks directly testing reward model performance. Examining their effect on policy model performance when used during PPO training, we find that these improved reward models have a large effect on GSM performance, but give marginal to no improvements across all other evaluations considered (§3).
@@ 3.1 Preference Data
Preference-based learning with existing datasets has the strongest effect on instruction following and truthfulness performance.
Our best models improve on the SFT model by over 8 points in these categories. In contrast, preference-based learning does not aid factuality, with all models remaining with 1 point of each other. This suggests that when using existing publically-available datasets, preference-based learning is most useful for improving chat-related abilities (instruction following, truthfulness) and learning stylistic features, but less strong at teaching new facts to a model. Interestingly, we observe that training on the Chatbot Arena data performs poorly on safety, suggesting Chatbot Arena volunteers generally prefer more toxic completions.
@@
Synthetic datasets generally outperform human-annotated and web-scraped datasets, especially in truthfulness. Additionally, using datasets collected by first getting per-aspect annotations (i.e., annotations from a human or model that score the helpfulness, harmlessness, etc. of the response independently) and then averaging across these scores tend to outperform datasets that rely only on overall judgements (i.e., just asking the annotator for an overall score instead of a per-aspect score). The two datasets that use this method, HelpSteer and UltraFeedback, display stronger or similar performance to datasets up to 15 times larger (e.g. HelpSteer vs HH-RLHF). We investigate the performance of varied sub-splits of UltraFeedback in App. E, which suggests that the use of per-aspect annotations is more important for performance than the quality of the models used to generate completions for the dataset.
@@
We now compare algorithms for learning from preferences, comparing the performance of DPO and PPO when the same base models and data are used (Table 2). We use exactly the same data for training DPO and PPO models, [^5] That is, we compare DPO models trained directly on each dataset with PPO models trained using reward models trained directly on that dataset and using prompts from the same dataset. Note that PPO uses additional generations from the model during training, but both approaches use the same amount of labelled data. and subsample larger datasets to 60,908 examples due to computational resources and only use these subsampled datasets during reward model and PPO training. For dataset choice, we use the top-performing dataset from each source type in Table 1 (StackExchange from Web, HH-RLHF from human, Ultrafeedback from synthetic). We also include the second-best performing dataset overall (Nectar) and an additional human-annotated dataset from a popular evaluation platform (Chatbot Arena 2023). Results in Table 2 indicate that:
@@
[…]
PPO improves on DPO in reasoning, coding and safety capabilities most.
@@ 3.3 Reward Models
Results in Table 3 indicate that either increasing the reward model dataset (‘Mix’) or reward model size (from 13B to 70B) improves direct RM performance. Surprisingly, we find that our 70B reward models perform best on BoN evaluation, while the 13B mixture model performs best on RewardBench. Both evaluations show that increasing the dataset mixture and increasing the dataset size can help, although increasing the dataset mixture is not as useful for further improving the 70B reward model. Although scaling model size helps with best-of-N, it does not improve RewardBench performance. Examining the per-split performance on RewardBench in App. H Table 13, the largest gaps between the 13B Mix RM and the 70B Mix RM is in reasoning (mostly code), suggesting that the larger model may not benefit much from the additional (somewhat noisy) data [^8] As the RewardBench reasoning subset is made up mainly of coding tasks, we hypothesise the presence of noisy StackExchange data may harm the 70B model, which has more capacity to fit to the larger mixture dataset than the 13B RM..
@@
We then test if our improved reward models lead to improved downstream policy models when used during PPO training. We perform PPO training using the UltraFeedback prompts during training and report our results on the right side of Table 3. Surprisingly, we find that improvements in reward models result in surprisingly small improvements in overall downstream performance. We see the largest (and only) overall improvement when using the 70B UltraFeedback RM, despite the fact that all improved RMs performed significantly better than the 13B UltraFeedback RM in RewardBench and Best-of-N evaluations. Additionally, the improvement from the 70B RM is largely driven by a large performance jump in GSM, as shown in Table 3, while other metrics stay largely similar. This suggests that it is difficult to translate improvements in reward models to the underlying policy. We note that most prior work examining reward models tends to examine either direct reward model performance [26, 61] or proxy reward [17, 40], rather than downstream performance. Our findings suggest that improving such evaluations may not necessarily translate to improved downstream performance. Additionally, we find that using the larger 70B UltraF. RM is less sensitive and performs well with lower KL penalty coefficient values than using the 13B UltraF. RM, and further examine the effect of the KL penalty coefficient on performance in App. I.
@@ 3.4 Policy Training Prompts
We first examine the effect of using policy training prompts targeted towards a particular downstream setting—specifically, math capabilities as evaluated by GSM. We construct three different prompt sets: prompts sourced from the GSM train set, math-related prompts from varied datasets, and 20K random prompts from UltraFeedback [^9] Note that this is only a third of all UltraFeedback data. We reduce the size to fairly compare to the small number of GSM8k prompts.. We further detail the approach used to identify math-related prompts in Appendix J. We then train models using PPO with each reward model from the previous section on each prompt set. Results in Fig. 3 demonstrate that:
@@
Surprisingly, we observe that using mixed prompts does not seem to improve performance in the generalist setting. When looking at code and math results specifically, we do not see consistent improvements using the altered prompt mixture, and our overall performance tends to drop. This is likely due to the already diverse nature of UltraFeedback, such that when looking at the whole dataset (i.e., not just the 20K subset in Figure 3), we are able to reach strong performance on math and coding evaluations. Altering the distribution away from other tasks slightly hurts the overall performance. We additionally found that training all the mined prompts in additional to all of UltraFeedback (i.e., not downsampling the combined prompt set) did not yield further improvements over the results shown in Table 4.
@@ 4 A Recipe for Learning from Preferences
Putting together all our findings from previous sections, we suggest a recipe for training a strong model using learning from preferences, as shown in Figure 1 and in Table 5. We take a high-quality, synthetic preference dataset, a large reward model, [^10] However, a smaller 13B RM performs almost as well if one is compute-constrained. and train it using PPO. If we additionally wish to focus on a specific domain, we can additionally collect domain-specific prompts for policy training. We find that our best model (Tülu 2+PPO 13B trained with the 70B UltraF. RM) outperforms our best DPO model and other popular models based off Llama 2 13B, including Llama 2 Chat, which has also undergone SFT and PPO training. Additionally, incorporating task-specific prompts into policy training may further improve performance when the prompts align closely with downstream evaluations, as shown in Figure 3. Finally, we also experiment with Llama 3.0 8B [35], finetuning on the Tülu2 Mix, and then training it using DPO and PPO with the same hyperparameters. We find that overall performance is significantly improved, and we observe similar trends as with our other experiments (DPO performing better than PPO, using a larger reward model improving performance, using mixed prompts not improving performance).
Model Factuality Reasoning Coding Truthfulness Safety Instr. Foll. Average Llama 2 13B Base 52.0 37.0 30.7 32.7 32.7 - - Llama 2 Chat 13B [52] 53.2 24.7 36.9 88.0 91.9 51.2 57.7 Nous Hermes 13B [51] 53.2 43.5 47.7 80.5 43.9 38.7 51.3 Vicuna 1.5 13B [64] 54.5 39.3 38.5 62.8 92.4 45.8 55.6 Tülu 2 13B SFT 55.4 47.8 45.1 56.6 91.8 44.2 56.8 Tülu 2+DPO 13B 55.3 50.9 45.9 69.3 91.9 52.8 61.0 Tülu 2+PPO 13B (13B UFRM) 56.0 52.0 47.7 71.5 91.8 54.4 62.2 Tülu 2+PPO 13B (70B UFRM) 55.4 53.9 47.3 72.3 91.9 55.8 62.8 Tülu 2+PPO 13B (70B UFRM+MP) 55.3 53.1 48.4 71.0 92.7 54.0 62.4 L3+Tülu 2 8B SFT 58.0 58.6 56.4 59.2 92.8 42.6 61.3 L3+Tülu 2+DPO 8B 59.4 56.2 55.6 71.4 91.7 50.4 64.1 L3+Tülu 2+PPO 8B (8B UFRM) 59.5 57.0 55.9 69.6 91.4 56.0 64.9 L3+Tülu 2+PPO 8B (70B UFRM) 58.5 60.8 55.0 72.8 91.8 55.8 65.8 L3+Tülu 2+PPO 8B (70B UFRM+MP) 58.3 40.6 48.2 62.4 91.0 53.0 58.9
Table 5: Putting together a recipe for preference-based learning: Performance of our best-performing models along with popular open models based on Llama 2 13B. ‘MP’ refers to using the mixed prompt set described in §4. ‘L3’ stands for experiments using Llama 3 as a base model. Using PPO with a large reward model performs best overall.
@@ 6 Conclusion
In this work, we have systematically explored the core components of learning from preference feedback and examined the relative impacts of each in turn on model performance across a wide range of evaluations. Our results suggest the following ordering of importance: preference data quality, algorithm choice, reward model quality, and finally targeted policy training prompts. Additionally, we find that using larger reward models can significantly improve math capabilities, but have marginal effects on other capabilities we evaluate in this work. Overall, we suggest a recipe for learning from preference feedback with currently available resources: best performance can be achieved by using a strong synthetic dataset (UltraFeedback), and using PPO training with a large reward model. Our work suggests that further exploring how to make better use of improved reward models is an important direction for further improving PPO-style approaches to learning from preference data. We plan to release models and code related to this paper and hope that our settings provide a firm base for future work further exploring learning from preferences for language models.
@@ Appendix A Limitations & Broader Impacts
As an empirical study with limited compute, our results are largely based on an in-depth examination over a single model suite (Tülu 2), using two base models (Llama 2 and 3). Additionally, our evaluation may not reflect all potential downstream use-cases - for example, we do not test multilingual performance. As such, examining how our results carry to multilingual settings or greatly differing base models is interesting future work. However, we have done our best to explore a wide variety of evaluations and datasets, using publicly available resources to aid in reproducibility. In terms of computational efficiency, we note that, while PPO does well, it comes with a significantly increased computational cost compared to DPO. We only consider performance in this study and don’t explicitly measure the relative computational cost of different methods (e.g., by measuring FLOPs cost of training with DPO vs PPO). There are also other aspects to keep improving PPO (e.g., doing new rounds of preference annotations to update the reward model for the changing distributions of the policy model), which we cannot exhaust in this paper, and we leave the exploration of them for future work.
@@ Appendix B Dataset Details
We additionally construct our ‘mix’ used for expanding the reward model training by a number of examples from each dataset as listed in Table 6. Overall, our ‘mix’ prefence set contains roughly 260k samples, over four times larger than just using UltraFeedback.
@@ Appendix E UltraFeedback Subset Study
# Samples Factuality Reasoning Coding Truthfulness Safety Inst. Foll. Average Llama 2 base - 52.0 37.0 30.7 32.7 32.7 - - Tulu 2 (SFT) - 55.4 47.8 45.1 56.6 91.8 44.2 56.8 UltraFeedback (FG; All) 60.908 55.3 50.9 45.9 69.3 91.9 52.8 61.0 Strong Model Gen.s 10,000 55.3 46.5 45.6 58.9 90.1 47.6 57.3 Middle Model Gen.s 10,000 55.2 48.5 44.5 60.6 91.5 48.2 58.1 Weak Model Gen.s 10,000 55.2 47.3 45.4 60.5 90.4 46.2 57.5 FalseQA Prompts 2,339 55.2 49.2 44.0 60.8 92.2 45.5 57.8 Evol Instruct Prompts 10,000 55.5 47.8 47.5 58.4 90.6 48.7 58.1 TruthfulQA Prompts 811 55.4 47.6 44.5 55.7 91.6 45.4 56.7 Flan V2 Prompts 2,0939 55.3 49.4 46.3 62.4 91.3 48.9 58.9 ShareGPT Prompts 19948 55.4 49.8 43.9 60.2 87.5 50.3 57.9 UltraChat Prompts 9929 55.4 50.6 47.5 59.9 89.9 48.1 58.6
Table 7: Performance of varied sub splits of UltraFeedback, considering both samples using generations from models of varying strength (weak, middle, strong model gen.) strong and only considering samples using prompts from single datasets (e.g., FalseQA, Flan V2). Sampling from different-quality models makes relatively little difference, while prompt choice matters. Additionally, different source datasets provide improvements in different evaluations.
@@ Appendix F Additional Details for PPO and DPO
We performed hyperparameter search for both DPO and PPO, sweeping across core hyperparameters in pilot experiments on HH-RLHF and UltraFeedback. We provide further details for both algorithms below.
@@ F.1 DPO
We testing multiple values for $\beta$ (0.1, 0.01, 0.001) and for learning rate (5e-6, 5e-7, 5e-8) in pilot experiments on HH-RLHF and UltraFeedback. Ultimately, we found that the hyperparameters suggested by Tunstall et al. [53] and followed by Ivison et al. [22] worked best (learning rate of 5e-7, $\beta$ of 0.01). This additionally involves 3 epochs of training with a learning rate of $5\times 10^{-7}$, with a linear warmup for 10% of training and linear cooldown for the duration. We use the open-sourced codebase used for training Tülu 2+DPO. [^17] https://github.com/hamishivi/EasyLM
@@ F.2 PPO
PPO formulates language generation as a Markov decision process (MDP), where each response $y$ is an episode, an action is a token $y_{t}$, and a state is the concatenation of a prompt and a partial response ($x\circ y_{<t}$). To construct token-level rewards $r_{t}$ that guide the RL training, the sequence-level reward $r$ is applied only to the last token, and each token is penalized by a KL term $-\beta\big{(}\log\pi_{\theta}(y_{t}\mid x\circ y_{<t})-\log\pi_{\text{ref}}(y_{t}\mid x\circ y_{<t})\big{)}$. Note that $\pi_{\theta}$ is the model we are training and $\pi_{\text{ref}}$ is a reference model (in our case, the initial SFT model we start training from). Formally, the token-level reward $r_{t}$ for each response token $y_{t}$ is defined as
@@
The prompt and continuation each has at most 1024 tokens. We use a batch size of 64 and the same for minibatch size, and train for 1 epoch over the prompt dataset. On each batch of prompts and rollouts, we train for 1 inner epoch. We use a sampling temperature $\tau=0.7$ for rollout, and a fixed KL penalty coefficient $\beta=0.05$. We optimize with AdamW with learning rate $\eta=1\times 10^{-6}$ with a linear warmup for 10% of iterations. We provide additional details on the hyperparameters used for PPO in Table 9. In experiments, we found that training with larger (70B) reward models benefited from a reduced KL penalty coefficient $\beta$, but using smaller reward models did not benefit (see App I for further discussion). As such, for runs using larger reward models, we use a KL penalty of $\beta=0.0325$ instead of $\beta=0.05$. This is in line with prior work suggesting that improved reward models are more difficult to overoptimize against, and so can benefit from lower KL penalties [17, 40]. We found these hyperparameters through experimenting in pilot experiments over HH-RLHF and UltraFeedback. We experimented with KL penalty coefficients of 0.01, 0.025, 0.0325, 0.05, as well as varying the batch size (including taking more samples per gradient step, and taking multiple steps over the same generations). We additionally explored training for up to 3 epochs and found that training beyond 1 epoch did not yield improved performance and occasionally resulted in the training collapsing entirely, and as such stuck to training for only 1 epoch.